Find the slope of a line perpendicular to the line that passes through the following points.
( 4, 11) and ( -2, 1)

Answer by Cromlix(4381)

Using gradient formula:
y2 - y1/ x2 - x1
1 - 11/-2 - 4
-10/-6
= 5/3
Lines that are perpendicular to
each other have slopes that multiply
together to give -1
m1 * m2 = -1
m1 = 5/3
m2 must equal -3/5
5/3 * -3/5 = -1
Slope = -3/5
Hope this helps.
:-)